# Zero-shot Generalizable Incremental Learning for Vision-Language Object Detection

### v1

@@ 4.1 Setup
Table 2: Main ablation results.
4.2 Comparison with Existing Methods
Experiments are conducted under both few-shot and full-shot settings. Since CL-DETR and OW-DETR are not designed for few-shot IOD, we only compare them under the full-shot setting. The results presented in Tab. 1 demonstrate that ZiRa consistently outperforms existing IOD methods in terms of “Avg” on downstream tasks. Moreover, ZiRa exhibits remarkable performance in preserving the zero-shot generalization ability of VLODMs under both the few-shot and full-shot settings. In particular, ZiRa surpasses CL-DETR and iDETR by substantial margins, with improvements of 13.91 and 8.71 AP under the full-shot setting in terms of “ZCOCO”, respectively. These results emphasize ZiRa’s effectiveness in maintaining the zero-shot AP and its potential as a superior solution for IOD. Distinctively, ZiRa, unlike CL-DETR and OW-DETR, requires minimal extra memory for a few branches without storing image exemplars and the model copy, making it more memory-efficient than other IOD methods.
4.3 Ablation Study
@@ V L ZiRa ZCOCO Avg hAP
Table 3: Comparison of learning on different modalities.
Branch Structure ZCOCO Avg hAP
@@
RDB (DB + Rep+) 46.06 59.73 52.01
Table 4: Comparison of different additional branch structures.
Branch Structure. To study the impact of the additional branch structure introduced in VOLDM, we compared the results of introducing a single branch (denoted as SB), dual branches (denoted as DB), and RDB. Besides RDB, both the output norm of SB and DB are also penalized with ZiL. As we can see in Tab. 4, the SB structure, due to its inability to introduce inter-branch division of labor, falls short in its ability to prevent forgetting in downstream tasks compared to the RDB. Moreover, we can see that merely introducing DB with ZiL cannot prevent forgetting on downstream tasks, on the contrary, it brings worse forgetting due to increased plasticity. RDB not only incorporates the mechanism of branch labor division but also fully utilizes $l_{hlrb}$ to learn to minimize the interference of HLRB, thereby more effectively preventing the forgetting on downstream tasks.
@@
Table 5: Results with varying $\lambda$ for ZiL.
Effect of $\lambda$. The hyperparameter $\lambda$ modulates the influence of ZiL within the overall loss function. In this study, we investigate the impact of varying values of $\lambda$ on model performance, as presented in Tab. 5. The results show that a small $\lambda$ value (e.g., $0.01$) tends to cause the model to overfit new tasks, resulting in reduced performance in the general domain. Conversely, a large $\lambda$ value (e.g., $1.0$) overly constrains the model, hindering its adaptation to downstream tasks. The optimal $\lambda$ value falls between these extremes, with the highest hAP observed when $\lambda=0.05$, which strikes a balance between “ZCOCO” and “Avg”.
Differentiated Learning Rate. We explore the effects of hyperparameter $\eta$. The results are detailed in Tab. 6. On the one hand, setting the LLRB’s learning rate as zero equivalently freezes it. Although this results in the highest “Avg” by preserving downstream task knowledge, the “ZCOCO” performance suffers considerably. This degradation is because HLRB, which has a limited parameter size, endures all the optimization pressure from $L_{zil}$ and $L_{cls}+L_{loc}$. Despite HLRB’s attempts to find a harmonious optimization path, it tends to emphasize downstream task learning over maintaining initial zero-shot generalizability, especially when the predominant loss $L_{cls}+L_{loc}$ outweighs $L_{zil}$. On the other hand, equating the LLRB’s learning rate to the HLRB’s is equivalent to creating two HLRBs inside the RDB. The enhanced plasticity that comes from the two-HLRB structure makes it challenging for ZiL to prevent it from gravitating toward directions detrimental to zero-shot generalizability, leading to diminished “ZCOCO” scores. Moreover, a significantly high LLRB learning rate can also precipitate a swift decline in downstream task knowledge retention. Optimal $\eta$ lies between these two cases, like 0.1 or 0.2, striking a balance between efficient new task adaptation and zero-shot generalizability protection.
@@
Table 6: Results with varying $\eta$.
5 Conclusion

### v3

@@ V L ZiRa ZCOCO Avg hAP
Table 5: Comparison of different additional branch structures. We compared the results of introducing a single branch (denoted as SB), dual branches (denoted as DB), and RDB (DB + Rep+).
Structure ZCOCO Avg hAP
@@
RDB 46.06 59.73 52.01
Branch Structure. To study the impact of the additional branch structure introduced in VOLDM, we compared the results of introducing a single branch (denoted as SB), dual branches (denoted as DB), and RDB (DB + Rep+). Besides RDB, both the output norm of SB and DB are also penalized with ZiL. As we can see in Tab. 5, the SB structure, due to its inability to introduce inter-branch division of labor, falls short in its ability to prevent forgetting in downstream tasks compared to the RDB. Moreover, we can see that merely introducing DB with ZiL cannot prevent forgetting on downstream tasks, on the contrary, it brings worse forgetting due to increased plasticity. RDB not only incorporates the mechanism of branch labor division but also fully utilizes $l_{hlrb}$ to learn to minimize the interference of HLRB, thereby more effectively preventing the forgetting on downstream tasks.
5 Conclusion
Limitations. While ZiRa has demonstrated its effectiveness in IVLOD tasks, there are still some limitations. First, the current implementation of ZiRa is based on the DETR architecture, which may not be optimal for all VLODMs. Second, the current ZiRa method is designed for the IVLOD task, and its generalization to other incremental learning tasks remains to be explored.
Acknowledgments
We would like to express our deepest gratitude to the anonymous reviewers and the community for their valuable feedback, which helped us enhance the quality of this work. Additionally, we are grateful to our colleagues for their insightful discussions and support throughout the development of this project.
Appendix
Effect of Hyperparameter
Impact of $\lambda$. The hyperparameter $\lambda$ modulates the influence of ZiL within the overall loss function. In this study, we investigate the impact of varying values of $\lambda$ on model performance, as presented in Tab. 6. The results show that a small $\lambda$ value (e.g., $0.01$) tends to cause the model to overfit new tasks, resulting in reduced performance in the general domain. Conversely, a large $\lambda$ value (e.g., $1.0$) overly constrains the model, hindering its adaptation to downstream tasks. The optimal $\lambda$ value falls between these extremes, with the highest hAP observed when $\lambda=0.05$, which strikes a balance between “ZCOCO” and “Avg”.
Table 6: Results with varying $\lambda$ for ZiL.
@@ Acknowledgments
Differentiated Learning Rate. We explore the effects of hyperparameter $\eta$. The results are detailed in Tab. 7. On the one hand, setting the LLRB’s learning rate as zero equivalently freezes it. Although this results in the highest “Avg” by preserving downstream task knowledge, the “ZCOCO" performance suffers considerably. This degradation is because HLRB, which has a limited parameter size, endures all the optimization pressure from $L_{zil}$ and $L_{cls}+L_{loc}$. Despite HLRB’s attempts to find a harmonious optimization path, it tends to emphasize downstream task learning over maintaining initial zero-shot generalizability, especially when the predominant loss $L_{cls}+L_{loc}$ outweighs $L_{zil}$. On the other hand, equating the LLRB’s learning rate to the HLRB’s is equivalent to creating two HLRBs inside the RDB. The enhanced plasticity that comes from the two-HLRB structure makes it challenging for ZiL to prevent it from gravitating toward directions detrimental to zero-shot generalizability, leading to diminished “ZCOCO" scores. Moreover, a significantly high LLRB learning rate can also precipitate a swift decline in downstream task knowledge retention. Optimal $\eta$ lies between these two cases, like 0.1 or 0.2, striking a balance between efficient new task adaptation and zero-shot generalizability protection.
Table 7: Results with varying $\eta$.
@@
Impact of Different Initial Values of the Scaling Factor. The scaling factor $s$ is a learnable parameter that affects the IVLOD’s performance. Its initial value could shape the model’s efficiency and adaptation capabilities. As documented in Tab. 8, contrary to initial expectations, a balanced scaling factor for both language and vision components, set at 1.00, does not lead to the best outcomes. Instead, an asymmetrical approach where the vision component’s scaling factor is reduced to 0.10 while maintaining the language component’s scaling factor at 1.00, or vice versa, appears to enhance overall performance metrics, including “ZCOCO", “Avg", and “hAP". Remarkably, the configuration where both language and vision scaling factors are set to 0.10 achieves superior results, yielding a slight improvement in “ZCOCO" and the highest “hAP" score among the tested scenarios. This suggests that a more conservative initial scaling factor may encourage the model to rely more on the base knowledge from pre-training, thus improving its generalization before adapting further through incremental learning.
Table 8: Impact of Different Initial Values of the Scaling Factor.
Language Vision ZCOCO Avg hAP
1.00 0.10 46.06 59.73 52.01
1.00 1.00 45.39 60.40 51.83
0.10 1.00 45.20 60.46 51.73
0.10 0.10 46.03 60.16 52.15
Impact of the Norm Types of ZiL
By default, we utilize the L1 norm to regulate ZiL. However, the choice of norm type for ZiL significantly impacts performance. As demonstrated in Tab. 9, L2 excels in ZCOCO, L1 strikes a balance between zero-shot and downstream performance, and Smooth L1 performs exceptionally well in downstream tasks while maintaining competitive zero-shot performance, even outperforming the L1 norm. Regardless of the norm type used, ZiRa remains effective, showcasing its versatility across different norm types.
Table 9: Results with Varying Norm Types of ZiL.
Norm Type ZCOCO Avg hAP
L1 46.06 59.73 52.01
L2 46.26 57.32 51.20
Smooth L1 46.10 60.07 52.17
Visualization
To more vividly showcase the effectiveness of the proposed ZiRa method, we also performed a visualization experiment. As shown in Fig. 6, using the pre-trained Grounding DINO directly to leverage its zero-shot detection capabilities allows for detecting many unseen objects, but fails to detect some. Upon incrementally fine-tuning with iDETR using training samples (with five classes per phase and one sample per class), the model will be able to recognize objects that are undetected by zero-shot detection. However, some objects that can be initially recognized by zero-shot detection are no longer detected. In contrast, after incremental fine-tuning with ZiRa, the images reveal that both familiar and novel objects are detected. This demonstrates ZiRa’s dual ability to prevent forgetting both pre-trained knowledge and information from downstream tasks.
Figure 6: Visualization results of both seen categories and unseen categories with Grounding DINO-based ZiRa.